A girl scout troop sold cookies. If the girls sold 5 more boxes the second week than they did the first, and if they doubled the sales of the second week for the third week to sell a total fo 431 boxes of cookies, how many did they sell each week? Please show me how to set up this problem. Thank you very much.
Cheyenne

Answer by scott8148(6628)

let x="first week sales"

"the girls sold 5 more boxes the second week" ___ x + 5

"they doubled the sales of the second week for the third week" ___ 2(x + 5) = 2x + 10

"a total fo 431 boxes of cookies" ___ x + x + 5 + 2x + 10 = 431

4x + 15 = 431